Determine the end (long run) behavior for the function below:

f(x)=x4−3x3+x−1
1
Expert's answer
2019-09-23T09:37:36-0400

The a long run behaviour of polynomial function is determined by its leading term (with the highest power). It means, that for long run asymptotics function "f(x) = {x^4} - 3{x^3} + x - 1" is equivalent to the simple function "g(x) = {x^4}".

"g(x)" is even function, so asymptotics for "f(x)" in both directions are the same. To be more exact, for

"x \to + \infty" we've got "f(x) \to + \infty"

and for

"x \to - \infty" we've got "f(x) \to + \infty" .

The graph of the given function "f(x)" is drawn below